# Knowledge Graph Large Language Model (KG-LLM) for Link Prediction

### 13 Mar 2024 v2

@@ Abstract
The task of predicting multiple links within knowledge graphs (KGs) stands as a challenge in the field of knowledge graph analysis, a challenge increasingly resolvable due to advancements in natural language processing (NLP) and KG embedding techniques. This paper introduces a novel methodology, the Knowledge Graph Large Language Model Framework (KG-LLM), which leverages pivotal NLP paradigms, including chain-of-thought (CoT) prompting and in-context learning (ICL), to enhance multi-hop link prediction in KGs. By converting the KG to a CoT prompt, our framework is designed to discern and learn the latent representations of entities and their interrelations. To show the efficacy of the KG-LLM Framework, we fine-tune three leading Large Language Models (LLMs) within this framework, employing both non-ICL and ICL tasks for a comprehensive evaluation. Further, we explore the framework’s potential to provide LLMs with zero-shot capabilities for handling previously unseen prompts. Our experimental findings discover that integrating ICL and CoT not only augments the performance of our approach but also significantly boosts the models’ generalization capacity, thereby ensuring more precise predictions in unfamiliar scenarios.
@@ 1 Introduction
  • Our analysis on real-world datasets confirms that our framework improves generative multi-hop link prediction in KGs, underscoring the benefits of incorporating CoT and IFT during training on both ICL and non-ICL evaluations.
  • By reorienting multi-hop link prediction from a traditional classification paradigm to a generative model, our framework significantly simplifies debugging for users.
@@ 2 Related Work
Recent development of large-scale language models, such as BERT [4], GPT [13], LLaMA [19], Gemini [18] and T5 [14], has significantly advanced NLP in many ways. These models are trained on vast amounts of text data and have demonstrated impressive performance on a wide range of NLP tasks. For example, Flan-T5 (Fine-tuned Language Net Text-To-Text Transfer Transformer) [25], as a variation of T5, is a Transformer-based language model that is pre-trained on a large corpus of diverse text data and fine-tuned on various downstream tasks. The text-to-text training approach makes LLMs particularly suitable for our generative link prediction task, where the model needs to generate the connection between multiple entities based on the input text. Besides, pre-training on large-scale data makes it possible for LLMs to generalize to unseen tasks or prompts that were not part of its training data [26]. In our study, we will utilize an unseen prompt “What is the relation between Node_id and Node_id?” instead of the original prompt “Is Node_id connected with Node_id?” to evaluate the model’s ability to generalize to unseen prompts.
@@
Figure 2: Models processed through the standard framework will be trained using the standard knowledge prompt dataset (left), whereas models processed via the KG-LLM framework will be trained on the KG-LLM knowledge prompt (right).
@@ 3.2 Knowledge Prompt
The knowledge prompt refers to a specialized prompt designed for KGs that enables the conversion of a given sequence of observed triples $P_{obs}$ into natural language, thereby facilitating the model learn better. By leveraging the knowledge prompt in the training process, the model is able to more effectively understand the underlying relationships and patterns present within KGs, thus improving overall performance in multi-hop link prediction tasks. In this paper, we define the two types of knowledge prompts, standard knowledge prompt and KG-LLM knowledge prompt:
Multi-hop Link Prediction Standard Prompt:
###Input:
[…]
Multi-hop Relation Prediction Standard Prompt:
###Input:
[…]
Multi-hop Link Prediction KG-LLM Prompt:
@@ 1. Yes
Multi-hop Relation Prediction KG-LLM Prompt:
@@
The two types of prompts demonstrate distinct approaches to enhancing model performance in multi-hop prediction tasks. Standard knowledge prompt stands out for its clarity and simplicity, providing a deep comprehension of the KG and facilitating improved prediction accuracy. On the other hand, KG-LLM knowledge prompt adopts a structured format that includes instructions and inputs. This approach involves breaking down complex and lengthy inputs into manageable, concise processes. By organizing the information systematically, the likelihood of the model misinterpreting the queries is significantly reduced. Such a methodical strategy ensures the model stays highly aware of the Chain-of-Thought (CoT) process throughout its training, thereby supporting its ability to interpret and navigate the KG.
As illustrated in Figure 2, our experiments utilize the previously described prompts: the standard knowledge prompt and the KG-LLM knowledge prompt. The standard knowledge prompt pairs a question ($Q$) with a target answer ($A$), reflecting the conventional format employed in language models. Conversely, the KG-LLM prompt enriches $Q$ with the KG CoT prompt and embeds the CoT reasoning process within $A$. It further introduces an instruction within $Q$ that narrows down the model’s response options (yes/no), thereby guiding it towards more precise outcomes by constraining the range of possible answers.
@@ 3.3 KG-LLM Framework
Our complete KG-LLM Framework is illustrated in Figure 1. In the initial stage, the original KG is taken as input. The preprocess function randomly selects a unique path in the KG and converts it into the KG-LLM knowledge prompt. During the fine-tuning phase, three distinct LLMs are utilized: Flan-T5-Large, LlaMa2-7B, and Gemma-7B. Different fine-tuning techniques are applied for each model within our framework. A global fine-tuning strategy is employed on Flan-T5 to boost its performance. For LlaMa2 and Gemma, a 4-bit quantized LoRA (Low-Rank Adaptation) modification [7] is implemented. This technique strategically adjusts model parameters with minimal computational overhead, thereby optimizing their efficiency and effectiveness in handling complex prediction tasks. During the training process, we use the cross-entropy loss function $L$. It calculates the difference between the model’s predicted token probabilities and the actual token probabilities in the expected output sequence. In the following equation, $n$ represents the length of the expected output sequence, $x$ stands for the input instruction, and $y_{i}$ denotes the i-th token in the expected output sequence.
@@
To evaluate our KG-LLM Framework, we train each model twice. As illustrated in Figure 2, the initial training session employs standard knowledge prompt inputs to establish a baseline. Subsequently, we instruction fine-tune the original models using our specialized KG-LLM knowledge prompt inputs. As shown in section 3.2, KG-LLM instruction falls under the classification category. By listing all possible options in the instruction, LLMs can more effectively follow and generate a response from the given choices. For instruction details, please refer to Appendix.
Upon the completion of the training phases, we subject each model to two inference task sets, each comprising two sub-tests: non-In-Context Learning (non-ICL) and In-Context Learning (ICL). The primary set of inference tasks is centered around multi-hop link prediction. Conversely, the secondary set probes the models’ generalization ability in multi-hop relation prediction, particularly with previously unseen prompts. Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the standard and KG-LLM frameworks.
@@ 4 Experiments
In this section, we conduct experiments to evaluate the effectiveness of the proposed KG-LLM framework to answer the following several key research questions:
@@ 4.1 Experimental Setup
To assess the effectiveness of our KG-LLM framework, we employ models subjected to the Standard framework as baselines and compare their performance across all experiments. In the non-ICL multi-hop link prediction testing, we also compare our LLMs approach with traditional approaches such as TransE [2], Analogy [9], CompleX [20], DistMult [30], and RESCAL [12] to evaluate performance.
@@
Flan-T5-Large (standard) 0.63 0.67 0.60 0.66
LlaMa2-7B (standard) 0.74 0.72 0.71 0.73
Gemma-7B (standard) 0.76 0.73 0.72 0.71
@@
Flan-T5-Large (standard) 0.70 0.74 0.61 0.63
LlaMa2-7B (standard) 0.83 0.85 0.81 0.80
Gemma-7B (standard) 0.86 0.87 0.82 0.83
@@ 4.2 Multi-hop Link Prediction without In-Context Learning
This section delves into the comparative analysis of the traditional approaches, standard framework and KG-LLM framework in the context of non-In-Context Learning (non-ICL) testing scenarios. Table 2 illustrates the baseline performance compared to the results achieved by our innovative approach. Traditional approaches are shown in the top section of the table, LLMs baseline in the middle section, and our framework in the bottom section.
Answer to Q1: The comparative results demonstrate that across all models, the implementation of the KG-LLM framework surpasses the traditional approaches and baseline standard framework across both evaluated datasets. This improved performance can be attributed to the KG-LLM framework’s utilization of knowledge prompts. These prompts enable LLMs to take advantage of the intricate network of relationships between entities and their interconnections within the KG. Furthermore, the process of Instruction Fine-Tuning (IFT) precisely tailors the models, honing their focus on the limited options. The evidence presented here underscores the efficacy of our KG-LLM framework, enriched with CoT and IFT, indicating its potential to significantly advance the domain of multi-hop link prediction tasks in real-world applications.
It should be noted that the performance metrics of the LlaMa2 and Gemma models within the standard framework exhibit good performance. This observation suggests that even absent the enhancements introduced by the KG-LLM framework, models can still perform multi-hop link predictions. Nevertheless, the pivotal challenge is overcoming complex multi-hop KG problems. The KG-LLM framework emerges as a decisive factor in this area.
@@ 4.3 Multi-hop Link Prediction with In-Context Learning
To evaluate the influence of In-Context Learning (ICL) on models subjected to both the standard and the KG-LLM frameworks, exclude the traditional approach as it lacks ICL capability. We conducted a comparative analysis using the same LLMs and testing inputs as in the previous section. The key distinction in this evaluation was the incorporation of an ICL example prior to each original testing input. This addition was meticulously designed to ensure task fairness and unbiased assessment; the ICL example shown in the appendix, derived from the training dataset, were restricted to a complexity level of three. This constraint aimed to prevent providing additional knowledge through the ICL example while furnishing a contextually relevant example. By adhering to these principles, we ensured a robust and fair evaluation of model performance under ICL conditions.
Table 3 reveals a notable enhancement in the performance of models under the standard framework, with LlaMa2 and Gemma models achieving an F1 and AUC score exceeding 80%. Remarkably, the adoption of ICL within the KG-LLM framework resulted in a significant performance uplift; notably, the Gemma model achieved a staggering 98% F1 score on the first dataset, while LlaMa2 recorded a 96% F1 score on the second dataset.
An interesting observation emerged regarding the Flan-T5 model under the standard framework on the second dataset, where performance did not improve or even slightly declined post-ICL integration. This phenomenon could be attributed to the increased length and complexity of the testing prompts induced by the added ICL context. While the inclusion of an ICL example generally aids model understanding, in certain cases, it might inadvertently be perceived as noise or irrelevant, potentially affecting the performance of the standard framework Flan-T5.
Answer to Q2: The experimental results suggest that the deployment of ICL does not uniformly improve performance across all models within the standard framework. However, for models refined through the KG-LLM framework, the integration of ICL consistently facilitates performance improvements. This finding underscores the nuanced role of ICL in model optimization, highlighting how effective it is in improving model performance within KG-LLM framework.
@@
Flan-T5-Large (standard) 0.14 0.06
LlaMa2-7B (standard) 0.23 0.17
Gemma-7B (standard) 0.22 0.19
@@
Flan-T5-Large (standard) 0.35 0.24
LlaMa2-7B (standard) 0.56 0.48
Gemma-7B (standard) 0.58 0.52
@@ 4.4 Multi-hop Relation Prediction without In-Context Learning
An intriguing observation emerged from the evaluation, shown in Table 4: all frameworks showcased limited performance in this task. Notably, the KG-LLM framework exhibited marginally superior performance relative to its standard counterpart. Upon examining the predictive outcomes, we reveal a predominant inclination towards negative relation predictions over positive ones, where models generated a ‘no relation’ output for negative instances. This trend was consistent across models operating under both frameworks. The difference in performance between the two frameworks mainly came from the fact that the KG-LLM models were better at accurately identifying some positive relationships.
@@ 4.5 Multi-hop Relation Prediction with In-Context Learning
In addition to evaluating the performance of models on the relation prediction task without In-Context Learning (ICL), this study further explores the impact of incorporating ICL into the task. The ICL example is shown in the appendix. The comparative results, as illustrated in Table 5, reveal a significant improvement in the generalization abilities of the models both under the standard framework and the KG-LLM framework, in contrast to the findings presented in Table 4. In particular, the LlaMa2 and Gemma models, when processed through the KG-LLM framework and augmented with ICL, achieved an accuracy exceeding 70% across both datasets.
@@ 5 Conclusions and Future Work
Our study introduces the Knowledge Graph Large Language Model Framework (KG-LLM Framework) as a promising solution for multi-hop generative prediction tasks in knowledge graph analysis. By leveraging techniques such as Chain of Thought (CoT) and Instruction Fine-tuning (IFT), models processed by KG-LLM framework have greatly enhanced the accuracy of predictions in KG-related tasks. Overall, the standard framework models have displayed relatively lower performance in multi-hop link prediction tasks, and the introduction of in-context learning (ICL) has further hindered its performance. However, the KG-LLM framework models have demonstrated strong performance in multi-hop link prediction tasks, both with and without the utilization of ICL. Additionally, it exhibits a moderate level of ability when aided by ICL in relation prediction tasks that are not seen during model pre-training. These findings highlight the importance of incorporating CoT and IFT methodologies in enhancing the performance of models, while also emphasizing the need for further exploration and improvement in relation prediction tasks and the ability of handling unseen tasks in general.

### 18 Mar 2024 v3

@@ Abstract
The task of predicting multiple links within knowledge graphs (KGs) stands as a challenge in the field of knowledge graph analysis, a challenge increasingly resolvable due to advancements in natural language processing (NLP) and KG embedding techniques. This paper introduces a novel methodology, the Knowledge Graph Large Language Model Framework (KG-LLM), which leverages pivotal NLP paradigms. We use our method to convert structured knowledge graph data into natural language, and then use these natural language prompts to finetune large language models (LLMs) to enhance multi-hop link prediction in KGs. By converting the KG to natural language prompts, our framework is designed to discern and learn the latent representations of entities and their interrelations. To show the efficacy of the KG-LLM Framework, we fine-tune three leading LLMs within this framework, employing both non-ICL and ICL tasks for a comprehensive evaluation. Further, we explore the framework’s potential to provide LLMs with zero-shot capabilities for handling previously unseen prompts. Our experimental findings discover that the performance of our approach can significantly boost the models’ generalization capacity, thereby ensuring more precise predictions in unfamiliar scenarios.
@@ 1 Introduction
  • By converting knowledge graphs into CoT prompts, our framework allows LLMs to better understand and learn the latent representations of entities and their relationships within the knowledge graph.
  • Our analysis on real-world datasets confirms that our framework improves generative multi-hop link prediction in KGs, underscoring the benefits of incorporating CoT and instruction fine-tuning during training.
@@ 2 Related Work
Recent development of large-scale language models, such as BERT [4], GPT [14], LLaMA [21], Gemini [20], T5 [15], and Flan-T5 [28] has significantly advanced NLP in many ways. These models are trained on vast amounts of text data and have demonstrated impressive performance on a wide range of NLP tasks. The text-to-text training approach makes LLMs particularly suitable for our generative multi-hop link prediction task, where the model needs to generate the connection between multiple entities based on the input text. Recent studies, concurrent work such as GraphEdit [6], MuseGraph [18] and InstructGraph [24], have shown that natural language is effective for representing structural data for LLMs. Besides, training on large-scale data makes it possible for LLMs to generalize to unseen tasks or prompts that were not part of its training data [29]. In our study, we will utilize an unseen prompt “What is the relation between Node_id and Node_id?” instead of the original prompt “Is Node_id connected with Node_id?” to evaluate the model’s ability to generalize to unseen prompts.
@@
Figure 2: Models processed through the ablation framework will be trained using the KG-LLM ablation knowledge prompt dataset (left), whereas models processed via the KG-LLM framework will be trained on the KG-LLM knowledge prompt (right).
@@ 3.2 Knowledge Prompt
The knowledge prompt refers to a specialized prompt designed for KGs that enables the conversion of a given sequence of observed triples $P_{obs}$ into natural language, thereby facilitating the model learn better. By leveraging the knowledge prompt in the training process, the model is able to more effectively understand the underlying relationships and patterns present within KGs, thus improving overall performance in multi-hop link prediction tasks. In this paper, we define the two types of knowledge prompts, KG-LLM knowledge prompt and KG-LLM (ablation) knowledge prompt:
Multi-hop Link Prediction KG-LLM Knowledge Prompt:
@@ 1. Yes
Multi-hop Relation Prediction KG-LLM Knowledge Prompt:
@@
Multi-hop Link Prediction KG-LLM(ablation) Knowledge Prompt:
###Input:
[…]
Multi-hop Relation Prediction KG-LLM(ablation) Knowledge Prompt:
###Input:
[…]
The two types of prompts demonstrate distinct approaches to enhancing model performance in multi-hop prediction tasks. KG-LLM knowledge prompt adopts a structured format that includes instructions and inputs. This approach involves breaking down complex and lengthy inputs into manageable, concise processes. By organizing the information systematically, the likelihood of the model misinterpreting the queries is significantly reduced. Such a methodical strategy ensures the model stays highly aware of the Chain-of-Thought (CoT) process throughout its training, thereby supporting its ability to interpret and navigate the KG. On the other hand, we remove the instructional component in the KG-LLM (ablation) knowledge prompt and the CoT reasoning process from the expected output. This approach stands out for its clarity and simplicity, providing a good comprehension of the KG and facilitating improved prediction accuracy.
As illustrated in Figure 2, our experiments utilize the previously described prompts: the KG-LLM knowledge prompt and the KG-LLM (ablation) knowledge prompt. The KG-LLM (ablation) knowledge prompt pairs a question ($Q$) with a target answer ($A$), reflecting the conventional format employed in language models. Conversely, the KG-LLM prompt enriches $Q$ with the KG CoT prompt and embeds the CoT reasoning process within $A$. It further introduces an instruction within $Q$ that narrows down the model’s response options (yes/no), thereby guiding it towards more precise outcomes by constraining the range of possible answers.
@@ 3.3 KG-LLM Framework
Our complete KG-LLM Framework is illustrated in Figure 1. In the initial stage, the original KG is taken as input. The preprocess function randomly selects a unique path in the KG and converts it into the KG-LLM and KG-LLM (ablation) knowledge prompt. During the fine-tuning phase, three distinct LLMs are utilized: Flan-T5-Large, LlaMa2-7B, and Gemma-7B. Different fine-tuning techniques are applied for each model within our framework. A global fine-tuning strategy is employed on Flan-T5 to boost its performance. For LlaMa2 and Gemma, a 4-bit quantized LoRA (Low-Rank Adaptation) modification [8] is implemented. This technique strategically adjusts model parameters with minimal computational overhead, thereby optimizing their efficiency and effectiveness in handling complex prediction tasks. During the training process, we use the cross-entropy loss function $L$. It calculates the difference between the model’s predicted token probabilities and the actual token probabilities in the expected output sequence. In the following equation, $n$ represents the length of the expected output sequence, $x$ stands for the input instruction, and $y_{i}$ denotes the i-th token in the expected output sequence.
@@
To evaluate our KG-LLM Frameworks, we train each model twice. As illustrated in Figure 2, the initial training session employs KG-LLM (ablation) knowledge prompt inputs to establish a baseline. Subsequently, we instruction fine-tune the original models using KG-LLM knowledge prompt inputs. The KG-LLM instruction falls under the classification category. By listing all possible options in the instruction, LLMs can more effectively follow and generate a response from the given choices. For instruction details, please refer to Appendix.
Upon the completion of the training phases, we subject each model to two inference task sets, each comprising two sub-tests: non-In-Context Learning (non-ICL) and In-Context Learning (ICL). The primary set of inference tasks is centered around multi-hop link prediction. Conversely, the secondary set probes the models’ generalization ability in multi-hop relation prediction, particularly with previously unseen prompts. Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the KG-LLM (ablation) and KG-LLM frameworks.
@@ 4 Experiments
In this section, we conduct experiments to evaluate the effectiveness of the proposed KG-LLM frameworks to answer the following several key research questions:
@@ 4.1 Experimental Setup
To assess the effectiveness of our KG-LLM frameworks, we compare their performance across all experiments. In the non-ICL multi-hop link prediction testing, we also compare our LLMs approach with traditional approaches such as TransE [2], Analogy [10], CompleX [22], DistMult [33], and RESCAL [13] to evaluate performance.
@@
Flan-T5-Large (ablation) 0.63 0.67 0.60 0.66
LlaMa2-7B (ablation) 0.74 0.72 0.71 0.73
Gemma-7B (ablation) 0.76 0.73 0.72 0.71
@@
Flan-T5-Large (ablation) 0.70 0.74 0.61 0.63
LlaMa2-7B (ablation) 0.83 0.85 0.81 0.80
Gemma-7B (ablation) 0.86 0.87 0.82 0.83
@@ 4.2 Multi-hop Link Prediction without In-Context Learning
This section delves into the comparative analysis of the traditional approaches, ablation framework and KG-LLM framework in the context of non-In-Context Learning (non-ICL) testing scenarios. Table 2 illustrates the baseline performance compared to the results achieved by our innovative approach. Traditional approaches are shown in the top section of the table, ablation framework is in the middle section, and KG-LLM framework is in the bottom section.
Answer to Q1: The comparative results demonstrate that across all models, the implementation of the KG-LLM framework surpasses the traditional approaches and ablation framework across both evaluated datasets. This improved performance can be attributed to the KG-LLM framework’s utilization of knowledge prompts. These prompts enable LLMs to take advantage of the intricate network of relationships between entities and their interconnections within the KG. Furthermore, the process of Instruction Fine-Tuning (IFT) precisely tailors the models, honing their focus on the limited options. The evidence presented here underscores the efficacy of our KG-LLM framework, enriched with CoT and IFT, indicating its potential to significantly advance the domain of multi-hop link prediction tasks in real-world applications.
We can observe that the performance of traditional models is not ideal. This performance is likely attributable to the complexities associated with multi-hop link prediction. Unlike the straightforward task of predicting a direct link between two nodes, traditional methods now must consider all intermediate nodes to arrive at a conclusion. This additional complexity consequently diminishes their performance. It should be noted that the performance metrics of the LlaMa2 and Gemma models within the ablation framework exhibit good performance. This observation suggests that even absent the enhancements introduced by the KG-LLM framework, models can still perform multi-hop link predictions. Nevertheless, the pivotal challenge is overcoming complex multi-hop KG problems. The KG-LLM framework emerges as a decisive factor in this area.
@@ 4.3 Multi-hop Link Prediction with In-Context Learning
To evaluate the influence of In-Context Learning (ICL) on models subjected to both the ablation and the KG-LLM frameworks, exclude the traditional approach as it lacks ICL capability. We conducted a comparative analysis using the same LLMs and testing inputs as in the previous section. The key distinction in this evaluation was the incorporation of an ICL example prior to each original testing input. This addition was meticulously designed to ensure task fairness and unbiased assessment; the ICL example shown in the appendix, derived from the training dataset, were restricted to a complexity level of three. This constraint aimed to prevent providing additional knowledge through the ICL example while furnishing a contextually relevant example. By adhering to these principles, we ensured a robust and fair evaluation of model performance under ICL conditions.
Table 3 reveals a notable enhancement in the performance of models under the ablation framework, with LlaMa2 and Gemma models achieving an F1 and AUC score exceeding 80%. Remarkably, the adoption of ICL within the KG-LLM framework resulted in a significant performance uplift; notably, the Gemma model achieved a staggering 98% F1 score on the first dataset, while LlaMa2 recorded a 96% F1 score on the second dataset.
An interesting observation emerged regarding the Flan-T5 model under the ablation framework on the second dataset, where performance did not improve or even slightly declined post-ICL integration. This phenomenon could be attributed to the increased length and complexity of the testing prompts induced by the added ICL context. While the inclusion of an ICL example generally aids model understanding, in certain cases, it might inadvertently be perceived as noise or irrelevant, potentially affecting the performance of the ablation framework Flan-T5.
Answer to Q2: The experimental results suggest that the deployment of ICL does not uniformly improve performance across all models within the ablation framework. However, for models refined through the KG-LLM framework, the integration of ICL consistently facilitates performance improvements. This finding underscores the nuanced role of ICL in model optimization, highlighting how effective it is in improving model performance within KG-LLM framework.
@@
Flan-T5-Large (ablation) 0.14 0.06
LlaMa2-7B (ablation) 0.23 0.17
Gemma-7B (ablation) 0.22 0.19
@@
Flan-T5-Large (ablation) 0.35 0.24
LlaMa2-7B (ablation) 0.56 0.48
Gemma-7B (ablation) 0.58 0.52
@@ 4.4 Multi-hop Relation Prediction without In-Context Learning
An intriguing observation emerged from the evaluation, shown in Table 4: all frameworks showcased limited performance in this task. Notably, the KG-LLM framework exhibited marginally superior performance relative to its ablation counterpart. Upon examining the predictive outcomes, we reveal a predominant inclination towards negative relation predictions over positive ones, where models generated a ‘no relation’ output for negative instances. This trend was consistent across models operating under both frameworks. The difference in performance between the two frameworks mainly came from the fact that the KG-LLM models were better at accurately identifying some positive relationships.
@@ 4.5 Multi-hop Relation Prediction with In-Context Learning
In addition to evaluating the performance of models on the relation prediction task without In-Context Learning (ICL), this study further explores the impact of incorporating ICL into the task. The ICL example is shown in the appendix. The comparative results, as illustrated in Table 5, reveal a significant improvement in the generalization abilities of the models both under the ablation framework and the KG-LLM framework, in contrast to the findings presented in Table 4. In particular, the LlaMa2 and Gemma models, when processed through the KG-LLM framework and augmented with ICL, achieved an accuracy exceeding 70% across both datasets.
@@ 5 Conclusions and Future Work
Our study introduces the Knowledge Graph Large Language Model Framework (KG-LLM Framework) as a promising solution for multi-hop generative prediction tasks in knowledge graph analysis. By leveraging techniques such as Chain of Thought (CoT) and Instruction Fine-tuning (IFT), models processed by KG-LLM framework have greatly enhanced the accuracy of predictions in KG-related tasks. Overall, the ablation framework models have displayed relatively lower performance in multi-hop link prediction tasks, and the introduction of in-context learning (ICL) has further hindered its performance. However, the KG-LLM framework models have demonstrated strong performance in multi-hop link prediction tasks, both with and without the utilization of ICL. Additionally, it exhibits a moderate level of ability when aided by ICL in relation prediction tasks that are not seen during model pre-training. These findings highlight the importance of incorporating CoT and IFT methodologies in enhancing the performance of models, while also emphasizing the need for further exploration and improvement in relation prediction tasks and the ability of handling unseen tasks in general.